OpenStudy (anonymous):

\[integrate => \int\limits 3^{2x}dx\]

OpenStudy (apoorvk):

Ohkay, Mr. Paul McCartney, let me provide you with a standard formula for this, and see if you can work this out: \[\Large \int a^{px+q}.dx = \frac{a^{px+q}}{p(\ln a)}\] a,p and q are constants.

OpenStudy (anonymous):

2x=t 2dx=dt \[\int\limits \frac{3^{t}dt}{2}\] \[=\frac{1}{2} \frac{3^{t}}{\ln3}\]

OpenStudy (anonymous):

\[\frac{1}{2}*3^{2x}*\frac{1}{\ln3}\]

OpenStudy (anonymous):

|Thanks! :)